OpenStudy (sadaf):

when object thrown upward direction from height h with initial velocity Vo : using these question..v(t)=dy/dt=Vo-gt ....... y(t)=Vo-1/2gt^2 a) What is the maximum height reached by the object? b) How long the object stays in the air? c) With what velocity does the object hits the ground?

OpenStudy (shayaan_mustafa):

Hello Sadaf. I want to tell you the most important thing about physics. That whenever you solve any numerical so kindly draw a pic. It solves half of your problem. Did you visualize this problem?

OpenStudy (shayaan_mustafa):

Sadaf simplifies your equations V=Vo-gt and y=Vo-0.5gt^2

OpenStudy (sadaf):

yes i made a roughly sketch

OpenStudy (sadaf):

okay fine....i got these eq. from following eq. d^2 y / dt^2 = -g

OpenStudy (shayaan_mustafa):

oh My GOD so much time you take to write 4 words madam. show your sketch. please. so we cud help you.

OpenStudy (shayaan_mustafa):

y did you differentiate them?

OpenStudy (shayaan_mustafa):

i want that you learn something from this question therefore i am not solving it for you and just helping you. kindly don't be aggressive

OpenStudy (sadaf):

me also want to learn i dont mind

OpenStudy (sadaf):

i made sketch in copy

OpenStudy (shayaan_mustafa):

can you made it here?

OpenStudy (shayaan_mustafa):

sorry i mean draw it here. can you?

OpenStudy (sadaf):

i tried but could not make

OpenStudy (shayaan_mustafa):

ok w8. let us try to do something. use simplified equations which I have given you above.

OpenStudy (shayaan_mustafa):

according to question a person standing on a height say h1. and throw a ball in upward direction. let ball goes to the height h2. so total height will be h=h1+h2 agree?

OpenStudy (sadaf):

no ball is thrown up from ground

OpenStudy (sadaf):

where h=0

OpenStudy (shayaan_mustafa):

"when object thrown upward direction from height h" but these are you wordings. ok do you know urdu language?

OpenStudy (sadaf):

oh yes sorry this is the statement of question

OpenStudy (sadaf):

my friend said me that ball is thrown from ground thats y i am confused

OpenStudy (sadaf):

yes i know urdu lang.

OpenStudy (shayaan_mustafa):

no according to question it is from height "h" kindly check your question

OpenStudy (shayaan_mustafa):

han to madam apna question check kren. ap question smjye phle. or part(a) 2nd equation se solve hoskta hy.

OpenStudy (sadaf):

yes u r right Sir

OpenStudy (sadaf):

sir i checked...ball is thrown upward from height h.....not from ground sorry

OpenStudy (sadaf):

V^2 - U^2 = 2aS right ????

OpenStudy (shayaan_mustafa):

ok then collect data.

OpenStudy (sadaf):

a=g, S=h, so h=Vo^2/2g

OpenStudy (sadaf):

U is initial velocity .... am i right ?

OpenStudy (shayaan_mustafa):

yes right.. now move ahead.

OpenStudy (shayaan_mustafa):

@sadaf. Sunye. hmax=Vo^2*(sintheta)^2/2g

OpenStudy (shayaan_mustafa):

heena help her. i am not going to solve for her. kindly teach her.

OpenStudy (shayaan_mustafa):

what?? why you disagree?

OpenStudy (shayaan_mustafa):

\[h _{\max}=V _{0}^{2} \sin ^{2}(\theta)/2g\]

OpenStudy (shayaan_mustafa):

http://en.wikipedia.org/wiki/Projectile_motion

OpenStudy (anonymous):

sorry my fault

OpenStudy (shayaan_mustafa):

ok g. ok.

OpenStudy (anonymous):

plz shayaan help her pl

OpenStudy (sadaf):

sir ap mujay kiyun nahe bata saktay :(

OpenStudy (shayaan_mustafa):

lol. i am asking to you and you are to me.

OpenStudy (shayaan_mustafa):

sir.. i am just 18 years old boy not sir.. okk.. dont say it again.

OpenStudy (sadaf):

ok orry

OpenStudy (shayaan_mustafa):

acha ap data collect kren. or post kren. dekhen me ye phle hi padh chuka hn. apne ni padha hua. agr me solve kerdeta hn apk lye to ye ap apne lye apne future se cheat krengi. solve it by yourself thank you.

OpenStudy (sadaf):

sorry

OpenStudy (sadaf):

kindly aap solve kar dain plz mera exam hai

OpenStudy (shayaan_mustafa):

ok ok thk hy.. for (a) hmen max height find kerni hy to max height pr final velocity 0 hti hy. agree?

OpenStudy (sadaf):

g

OpenStudy (shayaan_mustafa):

so, Vf=0 g=9.8 and Vi=Vo put this in your data.

OpenStudy (sadaf):

ten mintes k bad meri light off ho jani hai so ap answer bata dena mein check kar lun ge.... or meray do swal or b hay mein post kar dun?

OpenStudy (shayaan_mustafa):

g ap post kerden. or me light meri b jane wali hy.

OpenStudy (shayaan_mustafa):

ab ap solve kerne den. w8 kren.

OpenStudy (anonymous):

a)max height =v^2/2g hmax=0.05Vo^2 b)v=u+at 0=Vo+gt -Vo/g=t

OpenStudy (anonymous):

 now u know time will be same use ur mind a lil